OpenStudy (anonymous):

Using L'Hospital's Rule find: the limit as θ->0 of θ/tan(θ) Please show steps.

OpenStudy (abb0t):

take the derivative of the top and derivative of the bottom and then take the limit. if you still get an indetermined solution, apply l'hopitals again. If you see that it's not taking you anywhere, then you might want to considere playing around with tangent. Remember: \(tan(x)=\frac{sin(x)}{cos(x)}\)

OpenStudy (anonymous):

How do I take the derivative of a theta?

terenzreignz (terenzreignz):

Verily... \[\Large \frac{d}{d\theta}\theta = 1\] ;)

OpenStudy (abb0t):

think of it like being "x" just diffretnt vairbale.

OpenStudy (anonymous):

applying the l hopital rule we get \[\frac{ 1 }{ (\sec(x))^{2} } = (\cos(x))^{2}\]

OpenStudy (anonymous):

awesome